OpenStudy (anonymous):

Write each complex number in strandard form. √-1

OpenStudy (radar):

i

OpenStudy (watchmath):

Ok, you keep asking the same type of problem. The trick is to replace the -1 by \(i^2\). On this problem \(\sqrt{-1}=\sqrt{i^2}=i\). On your previous problem \(\sqrt{-144}=\sqrt{144i^2}=\sqrt{(12i)^2}=12i\).